OpenStudy (anonymous):

sqaure root of 10xy divide by square root of 2x^3

OpenStudy (campbell_st):

it can be rewritten as \[\frac{\sqrt{10xy}}{\sqrt{2x^3}} = \sqrt{\frac{10xy}{2x^3}}\] cancel common factors to obtain your answer.

OpenStudy (anonymous):

so 5x^2y?

OpenStudy (campbell_st):

not quite its \[\sqrt{\frac{5y}{x^2}}\]

OpenStudy (anonymous):

thanks